Problem 2 Salaries for new assistant professors at State College are normally distributed with a mean of $80000 and a standard deviation of $7000 We randomly survey 20 teachers from that district, let X be the salary of a randomly selected teacher at State, and let X be the average salary of 20 randomly selected teachers bxpan a.(3%)X~ b, (3%)8~ (3%) Normal Norma, D Find the probability that an individual teacher earns more than $65,000. P(X> 65000) d. (3%) Find the probability that the average salary for the sample is more than $65,000. P> 65000) e. (396) Frd te 90th percertie forthe average teachers salary for samples of 20 teachers. PR <a) 0.9 thena-
Answer #1

(a)

X ~Normal (80000, 70002)

b:

Since it is given that X has normal distribution so Xbar will also have normal distribution with following parameters

mu_{ar{x}}=mu=80000

and'

σ 7000 1565.2476

Hence,

X ~ Normal(80000. 156524762)

c)

The z-score for X = 65000 is

65000 80000 7000

So the required probability is

P(X > 65000) P(z > _2.14) 1-P(z < _2.14) 0.9838

d)

The z-score for 65000 is

z=rac{65000-80000}{1565.2476}=-9.58

So the required probability is

P(z > 65000) = P(z >-9.58) = 1-P(z <-9.58) = 1.0000

e)

Here we need z-score that has 0.90 area to its left. The z-score 1.28 has 0.90 area to its left. The required a is

a-80000 1565.2476

a80000 1565.2476 1.28

a-82003.516928 82003.52